OpenStudy (anonymous):

how do you write this x>7 or x<0 into a interval notation?

OpenStudy (anonymous):

Okay, so let's start with x<0 How do you do that part?

OpenStudy (anonymous):

I am trying to get the solution set, so i can graph it

OpenStudy (anonymous):

Alright, what's the interval notation for x<0?

OpenStudy (anonymous):

This was my problem 2|2x-7|+4>18 and the answer I came up with was x>7 or x<0 and I was needing a little help to write them as a solution set

OpenStudy (anonymous):

Yep. So how you you write the x<0 part?

OpenStudy (anonymous):

lol I am not sure :)

OpenStudy (anonymous):

Okay, so, x can be less than zero. Can it include zero?

OpenStudy (anonymous):

would it (infinate,0)?

OpenStudy (anonymous):

Not quite. It would have to be negative infinite. Do you know why?

OpenStudy (anonymous):

no it can't include 0 because its not great than are equal to

OpenStudy (anonymous):

\[(-\infty,0) \cup \] Okay, so we have this so far.

OpenStudy (anonymous):

because x is less than 0 right?

OpenStudy (anonymous):

Which says -infinite to zero, but not zero. If it included zero it would be like: (-infinite, 0] with the bracket Okay, so lets do the other part.

OpenStudy (anonymous):

x>7 What would that look like?

OpenStudy (anonymous):

then we have x>7 which i think it would be (infinate,7) sorry i am not sure how to the infinate sign lol

OpenStudy (anonymous):

yep, but for this one, you do (7, infinite) since infinite goes beyond seven.

OpenStudy (anonymous):

Ok I see where your coming from thanks!!

OpenStudy (anonymous):

So yep, you've got it. Altogether it looks like this: \[(-\infty, 0) \cup (7, \infty)\]

OpenStudy (anonymous):

No worries mate.